Find the volume and surface area of a right-circular cylinder for which the base has a diameter measuring 14 feet and a height which measures 3 yards.
Answer by fractalier(2101)

For a right cylinder,
V = (pi)(r^2)(h) = (pi)(7^2)(9) = 441(pi) cubic feet
A = 2(pi)r^2 + 2(pi)rh = 2(pi)((7^2) + 2(pi)(7)(9) = 98(pi) + 126(pi) = 224(pi) sq feet